OpenStudy (anonymous):

A farmer has some goats and hens on his farm. They have a total of 99 heads and legs. There are twice a many hens as goats. How many hens are there?

OpenStudy (anonymous):

x=goats y=hens 5x+3y=99 There are twice as many hens as goats ratio 1 hen : 2 goats y=2x then insert into 5x+3y=99 5x+3(2x)=99 11x=99 x=9 goats y=2(9) y= 18 hens

OpenStudy (anonymous):

Thank you so much....

OpenStudy (anonymous):

:)